OpenStudy (anonymous):

If (t - 2)^2 = 0, What is the value of (t + 3) (t + 6) ? If possible please show work/explanation...

Parth (parthkohli):

First, you gotta solve for \(t\).\[\sqrt{(t - 2)^2} = \sqrt{0} \implies t - 2 = 0\implies t = 2 \]Does that help?

Parth (parthkohli):

Now just find \((2 + 3)(2 + 6)\).

Parth (parthkohli):

Hi, James :) I respect your intent to help the asker, but you are not allowed to post answers here. http://openstudy.com/code-of-conduct Hope you took that positively :D